# Programming with Pixels: Computer-Use Meets Software Engineering

### v1

@@ 3 Programming with Pixels (PwP)
Modern software engineering (SWE) requires using multiple programming languages, tools, and modalities. Furthermore, it relies on a wealth of tools that required tremendous human effort to create—from linters, to visual code debuggers, to project management tools. Motivated by these observations, we create Programming with Pixels, an IDE environment that satisfies two properties: (i) it is expressive, meaning that an agent can perform any task that is achievable through a sequence of primitive operations (e.g., typing or clicking) within an IDE; (ii) an agent has access to any tool implemented within the IDE since using a tool amounts to performing a sequence of primitive actions.
3.1 PwP environment
We represent the PwP environment as a partially observable Markov decision process (POMDP). We define the PwP POMDP $\langle S,A,O,T,R\rangle$ as:
  • $\mathbf{S}$ is the set of states describing the IDE and the operating system context, including open files, active editor panels, and cursor positions.
  • $\mathbf{A}$ is the action space, encompassing all possible keyboard and mouse events. The atomic actions in PwP are provided by the xdotool library (Sissel, ), which allows specifying all possible keyboard and mouse events in a simple syntax. The specific action space varies based on the agent setting, described in (§5).
  • $\mathbf{O}$ is the observation space. The observation space varies based on the agent setting, described in (§5).
  • $\mathbf{T}$ is the transition function. Actions like inserting a character typically lead to deterministic changes in the IDE state, whereas background processes can introduce stochasticity in timing and responses.
  • $\mathbf{R}$ is the reward function that measures performance on a given task. For instance, after the agent finishes editing code to fix a bug, the environment can run a test suite on the updated files to compute a reward.
Trajectories in PwP can thus resemble real-world development work: an agent can fix a bug in a repository, use a suggestion tool to help with writing code, or create documentation. The IDE and its operating system environment track changes, run tests and return reward signals.
Table 1: Comparison of PwP with existing environments across different dimensions. PwP uniquely combines comprehensive IDE tool support with full multimodal support, general action space, and execution-based evaluation, while maintaining software engineering specificity.
Multi- General Observation State Tools Execution-Based SWE
Environment modal Action Space Space Checkpointing Reward Specific
GAIA (Mialon et al., 2023) Text Limited
SWE-Bench (Jimenez et al., 2023) Text Limited
SWE-Bench-MM (Yang et al., 2024b) Text, Image Text Limited
OpenHands (Wang et al., 2024b) Text, Image Tool Output + Browser SWE
TheAgentCompany (Xu et al., 2024) Text, Image Tool Output + Browser SWE
WEBSHOP (Yao et al., 2023) Text Browser
WEBARENA (Zhou et al., 2024) Text Browser
VWEBARENA (Koh et al., 2024a) Text, Image Screen Browser
BrowserGym (Chezelles et al., 2024) Text, Image Tool Output + Browser Browser
OSWORLD (Xie et al., 2024) Text, Image Screen OS
WindowsAgentArena (Bonatti et al., 2024) Text, Image Screen OS
PwP (Ours) Text, Image, Video, Audio Screen All IDE Tools
3.2 Key Features of Programming with Pixels
Expressive observation and action space.
A typical approach to building agents is to engineer a set of high-level actions for operations like “open file” or “list symbols in file”, and then engineer an environment that supports each action (Xia et al., 2024; Yang et al., 2024b; Wang et al., 2024b). Furthermore, agents receive textual outputs that are manually reformatted for each action. The key difficulty is that such engineering does not scale to a large number of actions or to the full range of software engineering tasks, and the agent may be specialized to the observation and action space. In contrast, PwP preserves standard screen-based user interaction. The agent can navigate IDE menus visually, move the cursor, and press keys. This makes the environment expressive: an agent can achieve any task that can be achieved through primitive actions in an IDE.
Full Spectrum of Developer Tools.
A modern IDE offers debuggers, linters, version control, refactoring utilities, integrated terminals, and many extensions. In particular, PwP is developed on top of VSCode, which has a rich set of built-in functionalities and extensions. Implementing each of these into an agent or environment would require significant human effort. However, PwP provides these capabilities out-of-the-box within a single environment. Agents can set breakpoints, execute code, use language-specific extensions, review error messages, or run tests in a consistent interface.
Multimodality and Language Agnosticism.
Because the IDE supports numerous programming languages through extensions and built-in modules, PwP naturally covers tasks across Python, Java, JavaScript, Lean, and more without requiring separate integrations. For instance, agents can use the same debugger interface for all languages, or use pre-existing linters provided by IDE extensions. Beyond screenshots, the environment provides video streams and audio, though we leave exploring these for future work.
Rich Feedback and State Access.
PwP can evaluate performance immediately using testing frameworks or compilation checks. For instance, if the agent modifies a file, the IDE can automatically trigger a build, update diagnostics, or run tests, generating real-time feedback. In cases requiring deeper inspection—such as verifying that a bug is truly fixed—the environment can reveal file-system changes, process states, or test results.

### v2

@@ 3 Programming with Pixels (PwP)
Testing computer-use agents (CUAs) on software engineering (SWE) requires an environment that captures the full complexity of modern software engineering, which involves multiple programming languages, tools, and modalities. Furthermore, a fair evaluation must provide access to the wealth of tools that human developers use and specialized AI SWE agents have access to, such as linters, visual debuggers, and even project management tools. To enable such evaluation, we create Programming with Pixels, an IDE environment that satisfies these two requirements. First, it is expressive, meaning that an agent can perform any task that is achievable through a sequence of primitive operations (e.g., typing or clicking) within an IDE, which includes a wide range of software engineering activities. Second, an agent has access to any functionality implemented within the IDE, since using IDE functionality amounts to performing a sequence of primitive actions.
PwP environment.
We represent the PwP environment as a partially observable Markov decision process (POMDP). We define the PwP POMDP $\langle S,A,O,T,R\rangle$ as follows. $\mathbf{S}$ is the set of states describing the IDE and the operating system (OS) context, including open files, active editor panels, and cursor positions. $\mathbf{A}$ is the action space, encompassing all possible keyboard and mouse events. The atomic actions in PwP are provided by the xdotool library (Sissel, ), which allows specifying all possible keyboard and mouse events in a simple syntax. The specific action space varies based on the agent setting, described in (§5). $\mathbf{O}$ is the observation space. The observation space varies based on the agent setting, described in (§5). $\mathbf{T}$ is the transition function. Actions like inserting a character typically lead to deterministic changes in the IDE state, whereas background processes can introduce stochasticity in timing and responses. $\mathbf{R}$ is the reward function that measures performance on a given task. For instance, after the agent finishes editing code to fix a bug, the environment can run a test suite on the updated files to compute a reward. Trajectories in PwP thus resemble real-world development work: an agent can fix a bug in a repository, use a suggestion tool to help with writing code, or create documentation. The IDE and OS environment track changes, run tests and return reward signals. In addition, we discuss five key features of PwP.
1. Expressive observation and action space. PwP provides computer-use agents with an unrestricted environment where they can attempt any software engineering task achievable through an IDE’s visual interface, as humans do. Unlike environments with predefined action sets, agents can navigate IDE menus visually, move cursors, and press keys to perform more complex actions.
2. Full Spectrum of Developer Tools. When evaluating computer-use agents on SWE tasks, it is imperative that they have a similar level of access to tools as specialized SWE agents, such as those with custom APIs for debuggers, linters, refactoring utilities, and more (Xia et al., 2024; Yang et al., 2024b). PwP provides all these tools through IDE’s visual interface, creating a comprehensive test of whether CUAs can leverage the same rich functionality that specialized agents access through APIs.
3. Multimodality and language agnosticism. CUAs promise generality across tasks and domains. Software engineering spans many languages such as Python, Java, JavaScript, Lean, and more, with tasks involving multiple modalities, such as text, images, data files, and PDFs, providing a rigorous test of this generality. In PwP, the same CUA must handle code generation, UI development, data science, and theorem proving without task-specific modifications. For agents requiring visual grounding support, we modified VSCode’s source code to provide rich DOM trees and Set-of-Marks annotations, ensuring fair evaluation across different CUA architectures.
4. Ease of verification. PwP provides direct access to the IDE’s internal state, file system, and OS processes for verification. When an agent modifies code, we can run test suites, check compilation, and verify correctness. This separation between agent interaction (visual) and evaluation (programmatic) makes it easier to verify task completion and provide other sources of feedback.
5. Future adaptability. Computer-use agents are improving rapidly, and so are software engineering agents. PwP is designed for future adaptability. First, adding new benchmarks is as simple as modifying configuration files. Second, PwP’s checkpointing is useful for search and RL training methods. Third, PwP’s gymnasium interface (Towers et al., 2024) provides a standard interface for evaluation and development. Finally, as agents improve and become capable of using more complex tools, the environment (IDE) would automatically incorporate these without architectural changes. This makes PwP an extensible platform for evaluating and developing computer-use agents.
Infrastructure and Implementation
PwP is deployed in a secure sandboxed docker environment, running open-source VSCode and a minimal operating system. Each container is isolated, preventing interference between experiments, ensuring parallel evaluation and facilitating reproducibility. We implement checkpointing for the environment state, which is especially useful for backtracking in search algorithms or training RL agents. The environment interfaces to VSCode using four channels for real-time screen capture, DoM information, and customizable configuration such as display, CPU/memory limits, etc. However, the complex interaction is abstracted away from the user, as they can simply interact with the environment through gymnasium python API (See Figure 9) and install the environment using a simple pip command. We refer to subsection A.3 for more details.
4 PwP-Bench
We introduce PwP-Bench, a benchmark containing 15 diverse software engineering tasks that span 14 programming languages and multiple modalities. Each task provides agents access to the IDE via the PwP environment. The goal of PwP-Bench is to test whether computer-use agents (CUAs) can handle the depth and breadth of software engineering activities.
Tasks.
PwP-Bench contains 5400 instances sourced from 13 existing code-generation datasets and 2 newly created by us. These tasks are designed to be representative of software engineering activities that take place within an IDE. Since the IDE is simply a computer program, in principle, these activities should be achievable by a general-purpose computer-use agent. We selected the tasks in PwP-Bench according to three key principles: (1) tasks must require substantial interaction with software engineering tooling, (2) each task should require multiple steps, and (3) the benchmark must cover multiple languages and modalities. Accordingly, tasks are grouped into four categories:
  • Code Generation and Editing: These tasks evaluate the ability to generate and edit code. This category includes datasets such as HumanEval for code completion, SWE-Bench (Jimenez et al., 2023) and SWE-Bench-Multilingual (Yang et al., 2025) for resolving pull requests, DSBench for data science tasks (Jing et al., 2024), and Res-Q (LaBash et al., 2024) or CanITEdit (Cassano et al., 2024) for code editing. Each dataset benefits from different IDE functionality. For example, SWE-Bench can take advantage of debuggers and linters, while DSBench may leverage an IPython kernel and extensions for analyzing large data files. Code editing tasks can leverage refactoring utilities and repository searches, covering varied input-output formats and end goals.
  • Multimodal Code Synthesis: These tasks involve creating code based on input images or other visual data. Examples include Design2Code (Si et al., 2024b) for UI development, Chart2Mimic (Shi et al., 2024) for generating Python code from chart images, SWE-Bench-MM (Yang et al., 2024b) for multimodal code editing, and DSBench tasks that rely on images, data files, or PDF documents for data analysis.
  • Domain-Specific Programming: These tasks focus on specialized fields such as ethical hacking (CTF) (Yang et al., 2023b) and interactive theorem proving (miniCTX) (Hu et al., 2024), which demand significant use and interaction with the IDE’s functionality. For example, theorem proving requires continuously inspecting goal states via the IDE, while CTF tasks involve analyzing images, running executables, or installing VSCode extensions (e.g., hexcode readers).
  • IDE-Specific and General SWE Tasks: Since code generation is only one aspect of software engineering, we introduce two novel task sets that evaluate broader SWE skills. The first, IDE Configuration, evaluates an agent’s ability to modify IDE settings such as themes, extensions, and preferences. These skills involve substantial interaction with the IDE, and are often a precondition for using IDE functionality such as new extensions. The second, which we term General-SWE, targets five different non-code activities: performance profiling, code refactoring, debugging bugs in standard libraries, UI mockup design, and code restoration. These tasks target practical software engineering tasks typically absent in conventional benchmarks. Full details are in Appendix B.2.
The distribution of tasks across categories and modalities is shown in Figure 10 in the Appendix. Computer-use agents that perform well across these tasks would demonstrate strong potential for automating diverse SWE activities across multiple languages, and working with varied input/output modalities such as text, images, data files, and other data types. Furthermore, taking advantage of the functionality provided by the agent’s environment is essential.
Benchmarking Design and Task Setup.
All tasks are evaluated within the PwP environment. Unlike traditional benchmarks, PwP-Bench presents agents with a realistic IDE environment: each agent receives an initial IDE state $S_{i}$ and an instruction $I$, with the goal to achieve a final state $S_{f}$ evaluated via execution-based criteria (e.g., unit tests). Among other capabilities, this setup tests whether CUAs can find relevant information from files, directories, and other resources, which is important for complex software development. Furthermore, a task is defined by a simple setup script that defines the initial IDE state, the instructions, and the evaluation logic. This makes it easy to add new tasks, allowing PwP-Bench to evolve as new benchmarks or better agents are developed.
PwP-Bench-Lite. Because PwP-Bench contains more than 5400 instances in total, running a full evaluation can be computationally expensive. To address this, we also provide PwP-Bench-Lite: a smaller subset of 300 instances. This subset preserves the overall difficulty and distribution while ensuring equal representation for each task, thereby making rapid experimentation more accessible.
5 Evaluating Agents in Programming with Pixels
We evaluate three distinct agent designs in the PwP environment to understand the capabilities and limitations of computer-use agents for software engineering tasks.
Computer-use agents. Computer-use agents interact with the IDE through primitive actions, i.e., keyboard and mouse inputs, while observing the interface visually through screenshots. Each agent operates in a turn-based manner, receiving a screenshot each turn and returning an action to progress toward the goal. Since most vision-language models without GUI-specific training struggle with raw pixel coordinates, we incorporate Set-of-Marks (SoM) (Yang et al., 2023a). With Set-of-Marks, an agent receives both the raw image and a parsed representation of available interface elements (e.g., buttons, text fields), allowing them to interact via element IDs rather than pixel coordinates. This design follows previous works (Xie et al., 2024; Koh et al., 2024).
Computer-use agents with File/Bash APIs. Computer-use agents are augmented with direct access to file-editing and bash commands through text APIs. The file-editing APIs include operations such as ‘read file’ and ‘string replace’, while bash operations allow command execution in the terminal. Agents receive screenshots only when requested via a screenshot action, rather than automatically each turn. This design strictly follows Anthropic’s computer-use implementation (Anthropic, 2024).
Specialized software engineering agents. For comparing how well current computer-use agents perform relative to specialized agents, we evaluate mini-sweagent (SWE-agent, 2024), an agent scaffold specifically designed for software engineering. Unlike computer-use agents that interact visually with the IDE, mini-sweagent operates entirely through text APIs. For multimodal tasks, it receives required images directly as input in its prompt. We chose mini-sweagent due to its near state-of-the-art performance on the widely-used benchmark SWE-Bench, as well as its flexibility for adapting to different programming tasks. See Appendix D.1 for implementation details.
Experimental setup.
We test multiple models as the parametrization for the two computer-use agent designs. Specifically, we test four vision-language models: Gemini-Flash-1.5, Gemini-Pro-1.5, GPT-4o, GPT-4o-mini, and we test five models with UI-specific training: closed-source Claude-3.5 Sonnet, Claude-3.7 Sonnet, Claude-4.0 Sonnet, and open-weights Qwen-2.5-VL and Qwen-GUI-Owl-32B. For the mini-sweagent, we test Claude-4.0 Sonnet. We keep the experimental setup consistent across all tasks and models: for each task instance, the maximum number of iterations is capped at 20 steps; if the agent either exhausts these steps or issues a stop command, the environment’s final state is evaluated using task-specific metrics (see Appendix B.2 for full details). For SWE-Bench related tasks, we further evaluate with a maximum of 250 steps in Appendix F. Due to computational and budget constraints, we evaluate on PwP-Bench-Lite, which has 300 task instances.
5.1 Results and Analysis
Table 1: Performance Evaluation of Different Agents on PwP-Bench by Task Categories. Best numbers are in bold, and best numbers for computer-use agents are underlined.
Model Code Generation Multimodal Domain-Specific General Overall
& Editing Code Generation Code Generation SWE Tasks Avg
Computer-Use Agents
Gemini-Flash 0.0% 4.3% 0.0% 0.0% 1.1%
GPT-4o-mini 0.8% 3.7% 0.0% 2.5% 1.7%
Qwen2.5-VL-72B 0.0% 4.3% 0.0% 5.0% 1.8%
GUI-Owl-32B 0.0% 0.0% 0.0% 22.5% 3.0%
Gemini-Pro 2.5% 5.7% 0.0% 7.5% 3.5%
GPT-4o 0.8% 12.4% 1.7% 10.0% 5.3%
Claude-Sonnet-3.5 10.7% 8.3% 5.0% 22.5% 10.5%
Claude-Sonnet-3.7 11.8% 28.5% 8.3% 27.5% 17.7%
Claude-Sonnet-4.0 16.0% 38.1% 6.7% 37.5% 22.9%
Computer-Use Agents with File/Bash APIs
Gemini-Flash 9.5% 11.7% 8.3% 2.5% 8.9%
GPT-4o-mini 23.6% 17.6% 15.0% 5.0% 17.8%
Qwen2.5-VL-72B 13.7% 11.8% 6.7% 7.5% 11.0%
Gemini-Pro 30.0% 16.7% 3.3% 12.5% 18.8%
GPT-4o 36.2% 41.9% 28.3% 10.0% 32.6%
Claude-Sonnet-3.5 47.9% 55.1% 43.3% 22.5% 45.5%
Claude-Sonnet-3.7 51.9% 58.7% 46.7% 27.5% 49.4%
Claude-Sonnet-4.0 53.4% 58.6% 43.3% 37.5% 50.7%
Software Engineering Agents
MiniSweAgent 49.4% 60.3% 40.0% 37.5% 48.8%
Table 1 summarizes performance across different agent architectures and base models over the four categories of PwP-Bench (task-wise results are in Table 6). As seen in the top half of the table, computer-use agents using only primitive keyboard and mouse actions achieve poor performance, with a maximum overall average of 22.9%. This is significantly lower than the software-engineering specific agent mini-sweagent, which achieves 48.8% accuracy. We attribute this poor performance primarily to limited visual grounding and an inability to interact effectively with the IDE, particularly for file editing and tool usage; see Section 5.1 for further analysis. Among all evaluated models, the Claude computer-use agent performs best, likely because it is specifically trained for UI interactions. We found that it can leverage basic IDE tools such as HTML live preview, chart visualization, and file navigation, boosting performance on tasks that require visual understanding and IDE navigation.
Nonetheless, when the same computer-use agents are granted access to just two text APIs (file editing and bash operations) we observe consistent improvements across all categories, with the maximum average accuracy reaching 50.7%, which, for some tasks is comparable to specialized state-of-the-art agent scores (see Appendix E). Interestingly, these operations could theoretically be performed without text APIs, since bash operations could be done through the IDE terminal and file editing through the file editor. While we show instances of CUAs attempting these visual operations in Section 5.1, they often make mistakes and are unable to recover from errors.
However, models still struggle to fully leverage the tooling available in the IDE. This is evidenced by poor performance on the ‘General SWE’ category, where tasks often require fewer than ten steps when using appropriate IDE tools. We analyze the poor performance on General SWE tasks further in the following sections, confirming that these tasks would become simpler if models could use IDE tooling more effectively. Overall, our results show that computer-use agents to have some facility for software engineering, but currently require better visual grounding, tool usage, and planning. In the following paragraphs, we analyze these strengths and deficiencies in more detail.
Claude Computer-Use Agent Demonstrates Basic IDE Tool Proficiency. Qualitatively, we found that Claude Computer-Use agent can use basic IDE functionalities, including file explorer navigation, file editing, search, browser-based live preview, and image generation and visualization capabilities. Figure 2 demonstrates the agent’s effective use of browser tools in UI replication tasks. Similarly, Figure 11 illustrates the agent’s ability to utilize multiple tools while editing specific lines in a repository, relying solely on screenshot observations and primitive keyboard/mouse actions.
Furthermore, we hypothesize that agents have additional latent abilities to use tools that can be activated through prompting or fine-tuning. To investigate this, we examined the project refactoring task (such as symbol renaming) in our ‘General-SWE’ benchmark, where Claude initially achieves 25% accuracy when attempting the task. However, when explicitly instructed to use precise tools (such as rename or move to file), its accuracy improves to 75% (see Appendix E).
Computer-Use Agents Demonstrate Poor Visual Grounding Capabilities.
Figure 4: Agent Hallucinating Screen Contents The agent correctly mentions, search results are displayed (green), it hallucinates an input field containing “disable import error” (red).
Figure 5: Agent Misidentifying UI Elements The agent fails to identify the correct input field, typing ‘50’ into the settings search bar instead of the word wrap column setting field (red arrow).
While, Claude Computer-Use agent is able to use basic IDE tools, we found that in general all current CUAs have significant limitations in visual grounding, i.e., the ability to understand the visual input and take actions on the visual IDE interface. We identify three primary failure modes. First, the agents can often fail to use the correct UI elements. For example, in Figure 5 the agent types in the search bar rather than the settings field, while in Figure 14 the model clicks the wrong location. Surprisingly, Set-of-Marks did not resolve these issues; agents would instead select incorrect elements.
Second, the agents often struggle to comprehend the current UI state, such as linter errors indicated by wavy underlines (Figure 12) or hallucinate screen contents (Figure 4). Finally, even when the agent can identify a simple error, such as incorrect indentation, it is often not able to fix the error due to struggling with clicking and typing in the proper locations. Furthermore, in Appendix F, models frequently completely ignore the visual state information and instead rely on completely memorized action sequences. Quantitatively, we found that 20% and 95% of trajectories have at least one visual grounding error in GPT-4o and Claude Sonnet-4.0, respectively (see Appendix F).
While grounding has been highlighted as a weak point of computer-use agents in web and OS domains (Koh et al., 2024; Xie et al., 2024), the limitations were primarily observed in models without UI-specific training. However, our work shows that even models explicitly trained for UI interaction, such as Claude Computer Use (Anthropic, 2024), exhibit these issues in PwP. We hypothesize that the deficiencies come from the IDE being particularly information-dense, as well as potentially not being covered by computer-use training datasets.
Agents Struggle to Use Advanced IDE Functionality. Although the best computer-use agent we tested could use basic IDE functionality, all agents lack the ability to leverage more sophisticated IDE tools. Specifically, we can see this through the low performance on the ‘General-SWE’ dataset, which focuses on software engineering activities (e.g., profiling, refactoring, debugging) that can be often completed without direct code edits. Although these tasks sometimes require only 4-5 steps when using appropriate IDE tools, agents achieve minimal performance, highlighting substantial room for improvement. Furthermore, we observed no successful uses of profilers, debuggers (even when explicitly instructed to) when performing the other tasks in our benchmark (see Appendix E).
Distribution of Functionality Used by Computer-Use Agents with File/Bash Operations.
Figure 6: Distribution of file/bash calls vs computer-use interaction for computer-use agents.
As we observed in Table 1, computer-use agents perform much better when they have access to file and bash API calls, which are based on text inputs and text outputs. A natural question is to what extent these are using the visual interface versus relying on text-only APIs. We study this in Figure 6, which shows the distribution of file/bash API calls versus computer-use interactions on four representative datasets. The figure shows a few interesting patterns. First, for HumanEval, agents rely entirely on file APIs. This is because HumanEval tasks involve simple function completions that are achievable without IDE interaction. The lower performance of pure CUAs on this task (25% compared to 100%) demonstrates their inability to perform basic file editing visually. Second, for SWE-Bench-MultiModal, surprisingly there are minimal computer-use interactions, primarily using screenshots to understand the open repository or occasionally attempting to open the built-in browser.
In contrast, the distribution shifts dramatically for Design2Code, where agents frequently open live preview tools to compare generated designs with reference images, and continuous refining the output (see Figure 2). In a similar vein, for VSCode tasks, the agents rely entirely on visual IDE functionality to update settings, install extensions, and edit themes. These patterns demonstrate that computer-use agents with file/bash APIs have some ability to choose between visual and API based interactions based on the task requirements. On datasets such as HumanEval, their performance improvements stem from bypassing their inability to visually perform edits, instead using text APIs.
Computer-Use Agents Are Rapidly Improving.
Figure 7: Performance of Claude-Sonnet Computer-Use Agents over time
Figure 7 compares the performance of Claude-Sonnet 3.5, 3.7, and 4.0 released over a period of 7 months. The line shows steady improvement in pure CUAs, with performance nearly doubling from 10.5% to 22.9%. Furthermore, from Table 1 we see that the gap between pure CUAs and CUAs with file/bash operations has steadily decreased from 35.0% to 27.8%. These results highlight that while a substantial gap remains, rapid progress is being made and continued improvements may eventually close this gap.
Leveraging the IDE functionality better would improve performance. While a single computer-use agent design can perform non-trivially across a wide variety of tasks, our analysis indicates that these models do not fully exploit domain-specific tools. To quantify the potential performance gains if agents could effectively use the IDE, we perform an “assisted” experiment. In this experiment, we manually engineered a set of IDE-based tool calls representing commonly used IDE functionalities (e.g., live HTML previews, repository structure, symbol outlines). Importantly, each API call is achievable using basic operations in the IDE, meaning that in principle, an agent could learn to perform it. See Appendix E for full details.
Figure 8: Assisted versus Computer-Use Agents
SWE-Bench Design2Code Chartmimic BIRD (T2 SQL)
Computer Use Agents 0% 23.5% 2.7% 0%
CUA + File/Bash 15% 48.1% 25.3% 7%
Assisted 19% 79.5% 61.6% 17%
Table 8 summarizes the performance improvements of assisted agents, highlighting an average gain of up to 13.3%. These results demonstrate that current CUAs have poor interaction capabilities with complex interfaces, yet there is significant scope for improvement. The results also suggest that in the near term, performance gains can be achieved by introducing specialized hand-engineered tools into computer-use agents and incorporating existing agent designs in our PwP environment.
6 Conclusion
We introduce Programming with Pixels, an environment designed to evaluate computer-use agents on software engineering tasks. We also introduce PwP-Bench, a diverse benchmark of 15 tasks spanning the breadth of software engineering across multiple languages and modalities. Our extensive evaluations of nine models reveal that pure computer-use agents relying solely on visual interaction perform poorly, while augmenting these agents with simple file and bash text APIs dramatically improves performance. Our analysis pinpoints poor visual grounding and an inability to leverage the rich set of functionality in the PwP environment as primary weaknesses. Despite these limitations, our findings show that CUAs are improving rapidly, signaling significant potential. PwP establishes software engineering as a natural domain for benchmarking whether generalist computer-use agents can reach specialist-level performance on sophisticated tasks.
7 Acknowledgements
We thank Daniel Fried, Graham Neubig, Zora Wang, Saujas Vaduguru, Atharva Naik, Riyaz Ahuja, and Weihua Du for helpful feedback. We thank Convergent Research and the OpenAI Researcher Access program. This work was supported in part by the National Science Foundation under Grant Nos. DMS-2434614 and DMS-2502281. Pranjal is supported by SoftBank Group–Arm Fellowship.
Appendix A Programming with Pixels (PwP) Environment
A.1 Tools
Previous methods have proposed use of various hand-engineered tools. For a fair comparison, all tools should be accessible in the PwP environment. Aas shown in Table 2, PwP natively supports all these tools.
Method Hand-engineered Tools Supported in PwP
CodeAct (Wang et al., 2024a) File Edit, IPython, Bash
SWE-agent (Yang et al., 2024a) Search File, Search Text, File Edit
A.2 Comparison with Other Environments
Table 3: Comparison of different environments across multiple dimensions
Computer-Use Execution-Based Specialized SWE
Environment Environment? Reward Domain Specific
GAIA (Mialon et al., 2023)
WEBSHOP (Yao et al., 2023)
WEBARENA (Zhou et al., 2024)
VWEBARENA (Koh et al., 2024)
BrowserGym (Chezelles et al., 2024)
OSWORLD (Xie et al., 2024)
AndroidWorld (Rawles et al., 2025)
WindowsAgentArena (Bonatti et al., 2024)
ScienceBoard* (Sun et al., 2025)
Cradle* (Tan et al., 2024)
PwP (Ours)
In Table 3, we compare PwP with existing environments across multiple dimensions. We evaluate environments along the following dimensions:
  • Computer-use environment: Whether the environment is designed for computer-use agents, and thereof whether it supports multimodal interaction.
  • Specialized: Whether the environment is designed for general and basic tasks, such as web navigation, or is it designed for a more sophisticated, specialized and potetntially economically important tasks. Only Cradle (Tan et al., 2024) and ScienceBoard (Sun et al., 2025) are specialized for Game Playing and using Scientific softwares respectively.
Further, ours support other engineering features that others do not. For instance, PwP also support streaming video and audio, something other environments do not support out of the box. Further, unlike environments such as OS-World, which require manual creation of environment image, PwP is natively docker based, and is based on simple scripts, that can be easily used to modify startup scripts and other configurations for future adaptations. Finally, we also specifically suppport state checkpointing which supports storing file system and complete process state, and is especially useful for search-based methods.